# Bounds on the number of squares in recurrence sequences

### v1

@@ 3.4. Miscellaneous Lemmas
From (3.3), along with $y_{k},y_{\ell}\geq 4b^{2}$, we have
@@
$\left|\omega^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|=\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
@@
for $\zeta_{4}^{\prime}\neq\zeta_{4}$. Since this exceeds $0.403$, it follows that $\zeta_{4}=\zeta_{4}^{(k)}/\zeta_{4}^{(\ell)}=\pm 1$, the last equality holding by our assumption in the statement of this lemma. ∎
In Lemma 3.9, we establish a gcd result for elements of a generalisation of our sequences. This will help us prove Lemmas 3.10 and 3.11. Lemma 3.11 is used in the proof of Proposition 4.1, as well as in the proof of Theorem 1.4. Lemma 3.11 will be particularly important for showing that the possible exceptions to Conjecture 1.2 in the statement of Theorem 1.4 can only possibly occur for $u=1$ or $u=2$.
@@
Let $\left(x_{k}\right)_{k=0}^{\infty}$ and $\left(y_{k}\right)_{k=0}^{\infty}$ be sequences defined by $x_{k}+y_{k}\sqrt{d}=\left(a+b\sqrt{d}\right)\varepsilon^{k}$, where $a,b,d$ are positive integers, $d$ is not a square and $\varepsilon\neq\pm 1$ is a unit in $\mathbb{Q}\left(\sqrt{d}\right)$.
@@
We write $\varepsilon=\left(t+u\sqrt{d}\right)/2$, where $t$ and $u$ are integers.
@@
We show that
@@
For $k\geq 3$, we use induction and the recurrence relations for $x_{k}$ and $y_{k}$:
@@
(3.34) $\left(\frac{u\left(3t^{2}+du^{2}\right)}{8N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{t\left(t^{2}+3du^{2}\right)}{8N_{\varepsilon}^{2}}\right)y_{k}$ $=-N_{\varepsilon}y_{k-3}\quad\text{ and }$
$\left(\frac{t\left(t^{2}+3du^{2}\right)}{8N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{du\left(3t^{2}+du^{2}\right)}{8N_{\varepsilon}^{2}}\right)y_{k}$ $=N_{\varepsilon}x_{k-3},$
@@
Since $3t^{2}+du^{2}=-N_{\varepsilon}+4t^{2}\equiv 4+4\equiv 0\pmod{8}$ and $t^{2}+3du^{2}=N_{\varepsilon}+4u^{2}\equiv 4+4\equiv 0\pmod{8}$, all the coefficients of $x_{k}$ and $y_{k}$ in the relationships in (3.34) are integers.
@@
Since $\left(\left(t+u\sqrt{d}\right)/2\right)^{2}=\left(t^{2}+u^{2}d\right)/4+(tu/2)\sqrt{d}=\left(t_{2}+u_{2}\sqrt{d}\right)/2$ with $t_{2}$ and $u_{2}$ both odd, we see that $x_{2}$ and $y_{2}$ are not integers either.
However, $\left(\left(t+u\sqrt{d}\right)/2\right)^{3}=\left(t^{3}+3tu^{2}d\right)/8+\left(\left(3t^{2}u+u^{3}d\right)/8\right)\sqrt{d}=\left(t_{3}+u_{3}\sqrt{d}\right)/2$ with both $t_{3}$ and $u_{3}$ even. Hence $x_{3}$ and $y_{3}$ are both integers.
By an inductive argument, one finds that $x_{k}$ and $y_{k}$ are both integers if and only if $3|k$, when $t$ and $u$ are both odd and $a$ and $b$ have opposite parities. So we consider only $\gcd\left(x_{3k},y_{3k}\right)$.
By (3.34), we have $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-3},y_{k-3}\right)|\cdots|\gcd\left(x_{0},y_{0}\right)=1$.
(ii-b) If $a$ and $b$ have the same parity, then $y_{1}=(au+bt)/2\in\mathbb{Z}$. Similarly, $x_{1}\in\mathbb{Z}$. Using the recurrence relations, we find that all the $x_{k}$’s and $y_{k}$’s are integers.
Using (3.34), we find that if $k\equiv k_{1}\pmod{3}$ with $0\leq k_{1}\leq 2$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k_{1}},y_{k_{1}}\right)$.
@@
Suppose first that $y_{0}$ is odd. Since $t$ is odd, we find that $y_{2}=ty_{1}+y_{0}$ is even if $y_{1}$ is odd and is odd otherwise.
In the first case (when $y_{1}$ is odd), we have $\gcd\left(x_{1},y_{1}\right)|\left(\gcd\left(x_{0},y_{0}\right)\right)$ and hence
@@
Now we suppose that $x_{0}$ is odd. Since $t$ is odd, we find that $x_{2}=tx_{1}+x_{0}$ is even if $x_{1}$ is odd and is odd otherwise. As above, we find that
@@
(iii) Now suppose that $t$ is even and $u$ is odd. In this case, $4|d$. In fact, if $4|t$, then $d\equiv 4\pmod{8}$ since $\varepsilon$ is a unit, and if $t\equiv 2\pmod{4}$, then $d\equiv 0\pmod{8}$.
Since $y_{1}=(au+bt)/2$, we see that $y_{1}\in\mathbb{Z}$ if and only if $a$ is even. Hence all the $y_{k}$’s are integers if $a$ is even. Since $t$ is even, using the recurrence relation, we see that for $a$ odd, $y_{k}\in\mathbb{Z}$ when $k$ is even and $y_{k}\not\in\mathbb{Z}$ otherwise.
Since $x_{1}=(at+bdu)/2$, we see that $x_{1}\in\mathbb{Z}$. Hence by the recurrence relation, $x_{k}\in\mathbb{Z}$ always holds.
As in case (ii) above, we use (3.34). If $k\equiv 0\pmod{3}$, then
$\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{0},y_{0}\right).$
If $k\equiv 1\pmod{3}$, then
$\gcd\left(x_{k},y_{k}\right)\text{ divides }\left\{\begin{array}[]{ll}\gcd\left(x_{1},y_{1}\right)&\text{if $a$ is even},\\ \gcd\left(x_{4},y_{4}\right)&\text{if $a$ is odd}.\end{array}\right.$
We need to stop at $\gcd\left(x_{4},y_{4}\right)$ when $a$ is odd because $y_{1}$ is not an integer in this case. By (3.32), we see that the lemma holds if $a$ is even. By (3.33) and observing that $2|t$ and $4|d$, we see that $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-2},y_{k-2}\right)$. So $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{0},y_{0}\right)$ when $a$ is odd.
If $k\equiv 2\pmod{3}$, then from (3.32) and (3.33), along with the observation when $k\equiv 1\pmod{3}$ and $a$ is odd, we have
$\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{2},y_{2}\right)|\gcd\left(x_{0},y_{0}\right).$
@@
Let the sequences $\left(x_{k}\right)_{k=-\infty}^{\infty}$ and $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be as defined in (1.1). Suppose that $k\neq 0$, $\gcd(a,b)=1$ and that $x_{k}$ and $y_{k}$ are both integers. Using the notation of Subsection 2.1 with $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$, $u_{1}=2x_{k}$ and $u_{2}=\pm 2\sqrt{N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)}$, then
@@
Put $v_{p}\left(g^{2}\right)=\ell\geq 0$. We will prove that $v_{p}\left(\gcd\left(a^{2},d\right)\right)=\ell$. We also put $v_{p}\left(g_{1}^{2}\right)=\ell_{1}$ and $v_{p}\left(g_{2}\right)=\ell_{2}$. Note that $\ell_{2}=0$ or $1$, since $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$ (recall from (2.4) that $g_{2}|t^{\prime}$ and $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$ here). From (2.4), we have $g_{3}|4$, so $v_{p}\left(g_{3}\right)=0$ for odd primes, $p$. Hence $\ell=\ell_{1}+\ell_{2}$.
@@
If $p\mid\operatorname{sf}\left(N_{\alpha}\right)$, then $p^{\ell_{1}+1}\parallel N_{\alpha}$. We must have $\ell_{2}=0$, since otherwise $\ell_{1}+1<\ell_{1}$ and we get a contradictions to $p^{\ell+1}|N_{\alpha}$. In this case, $p\nmid\left(x_{k}^{2}/g_{1}^{2}\right)$. So $p^{\ell_{1}}\parallel x_{k}^{2}$. But this contradicts $p^{\ell+1}|x_{k}^{2}$. Hence $p^{\ell+1}\nmid\gcd\left(a^{2},d\right)$. Therefore, for all primes $p>2$, $v_{p}\left(g^{2}/\gcd\left(a^{2},d\right)\right)=0$.
@@
We have $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}\right)\right)\geq\ell+2$. Also $g_{2}=1$. Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}/g_{3}\right)\geq\ell$, recalling that $g_{3}=1,2$ or $4$.
(2.2) Suppose that $2|\left(N_{\alpha}\right)$.
Again, we break this case into subcases.
(2.2.1) If $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=\ell$, then we saw above that $v_{2}\left(N_{\alpha}\right)\geq\ell+1$. So $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}/2\right)\right)\geq\ell+2$. Here $g_{2}\geq 1$, so as in case (2.1), we have $v_{2}\left(g^{2}\right)\geq\ell$, as required.
(2.2.2) If $v_{2}\left(a^{2}\right)\neq v_{2}\left(db^{4}\right)$, then $v_{2}\left(N_{\alpha}\right)=\ell$. So $v_{2}\left(g_{1}^{2}\right)=\ell+1$. Since $2|\operatorname{sf}\left(N_{\alpha}\right)$, it follows that $\ell$ must be odd.
From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$ and $v_{2}\left(N_{\alpha}\right)=\ell$, we have either $v_{2}\left(x_{k}^{2}\right)=\ell$ and $v_{2}\left(dy_{k}^{2}\right)\geq\ell+1$; or else $v_{2}\left(x_{k}^{2}\right)\geq\ell+1$ and $v_{2}\left(dy_{k}^{2}\right)=\ell$. The first case is not possible because $\ell$ is odd. In the latter case, we must have $v_{2}(d)=\ell$ and $2\nmid y_{k}$. Also, $v_{2}\left(u_{1}/g_{1}\right)=(1/2)v_{2}\left(4x_{k}^{2}/g_{1}^{2}\right)\geq(1/2)\left(\ell+3-(\ell+1)\right)=1$, so $v_{2}\left(g_{2}\right)=1$. Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}g_{2}/g_{3}\right)=v_{2}\left(g_{1}^{2}g_{2}/4\right)=\ell$. ∎
@@
To establish the equality in equation (3.35), we use Lemma 3.10. From that lemma, we can write $g^{2}=2^{m}\gcd\left(a^{2},db^{4}\right)$, where $m\geq 0$. So, squaring both sides of the equality in (3.35), applying the definitions of $b^{\prime}$ and $d^{\prime}$ and simplifying, it is equivalent to
@@
To show that this equality holds, we prove that if $v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)>6$, then $m=0$ (i.e., $g^{2}=\gcd\left(a^{2},db^{4}\right)$). Putting $\ell=v_{2}\left(\gcd\left(a^{2},db^{4}\right)\right)$, we must have $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=v_{2}\left(d\right)=\ell$. The second-last equality holds because $\gcd(a,b)=1$. From Lemma 3.9, we have $\gcd\left(x_{k},y_{k}\right)=1,2$. We will show that $\gcd\left(x_{k},y_{k}\right)=2$ is not possible here.
From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $\left(x_{k}/2^{\ell/2+1}\right)^{2}-\left(d/2^{\ell}\right)\left(y_{k}/2\right)^{2}N_{\alpha}/2^{\ell+2}$. Since $v_{2}\left(N_{\alpha}\right)\geq\ell+5$, it follows that $x_{k}/2^{\ell/2+1}$ and $y_{k}/2$ are both odd. Arguing mod $8$, we find that $d\equiv 1\pmod{8}$ and that $t$ and $u$ must both be even.
From (1.2), we can write $4y_{k}=\left(b^{2}\left(2du^{2}\pm 4\right)\pm 2atu\right)$ for some positive integers $t$ and $u$ with $t^{2}-du^{2}=\pm 4$. Since $t$ and $u$ are both even, it follows that $v_{2}\left(4y_{k}\right)=v_{2}\left(2du^{2}\pm 4\right)=2$.
If $a$ is even, then $b$ is odd. So if $u$ is also odd, then $4|(2atu)$, while $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=1$. Hence $2\nmid y_{k}$.
If $u$ is even, then $16|(2atu)$, while $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=2$. Hence $2\nmid y_{k}$.
We saw above that $\gcd\left(a^{2}\right)=\gcd\left(db^{4}\right)$, so if $a$ is odd, then so is $b$. Hence $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=2$, since $u$ is even. Here $8|(2atu)$, so $v_{2}\left(4y_{k}\right)=2$.
In none of these cases did we find that $2|y_{k}$. So
$v_{2}\left(g_{1}^{2}\right)=\min\left(v_{2}\left(4x_{k}^{2},4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=v_{2}\left(4x_{k}^{2}\right)=\ell+2.$
@@
Suppose that $b=1$, $k\neq 0$, that $x_{k}$ and $y_{k}$ are both integers and that $g$ and $\mathcal{N}_{d^{\prime},4}$ are as above.
@@
We have $|g|\mathcal{N}_{d^{\prime},4}\geq 2^{1+\min\left(v_{2}(N_{\alpha})/2,2\right)}\geq 2$ from (3.35). Since $b=1$, from Lemma 3.5(c), we find that $y_{k}>\left|N_{\alpha}\right|/4$ holds, so

### v2

@@ 3.4. Miscellaneous Lemmas
From (3.22), along with $y_{k},y_{\ell}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, we have
@@
$\left|\omega^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|$
$=$ $\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
@@
for any $4$-th root of unity, $\zeta_{4}^{\prime}$, with $\zeta_{4}^{\prime}\neq\zeta_{4}$. Since this exceeds $0.403$, it follows that $\zeta_{4}=\zeta_{4}^{(k)}/\zeta_{4}^{(\ell)}=\pm 1$, the last equality holding by our assumption in the statement of this lemma. ∎
In Lemma 3.10, we establish a gcd result for elements of a generalisation of our sequences. This will help us prove Lemmas 3.11 and 3.12. Lemma 3.12 is used in the proof of Proposition 4.1, as well as in the proof of Theorem 1.4. Lemma 3.12 will be particularly important for showing that the possible exceptions to Conjecture 1.2 in the statement of Theorem 1.4 can only possibly occur for $u=1$ or $u=2$.
@@
Let $\left(x_{k}\right)_{k=0}^{\infty}$ and $\left(y_{k}\right)_{k=0}^{\infty}$ be sequences defined by $x_{k}+y_{k}\sqrt{d}=\left(a+b\sqrt{d}\right)\varepsilon^{k}$, where $a,b,d$ are positive integers, $d$ is not a square and $\varepsilon=\left(t+u\sqrt{d}\right)/2\neq\pm 1$, with $t$ and $u$ integers, is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$.
@@
We start with some relationships that we will require for the proof itself.
Throughout the proof, we will use the fact that both the $x_{k}$’s and $y_{k}$’s satisfy the recurrence relation
(3.32) $u_{k}=\operatorname{Tr}\left(\varepsilon\right)u_{k-1}-N_{\varepsilon}u_{k-2}=tu_{k-1}\pm u_{k-2}.$
Next, we show that
@@
For $k\geq 3$, we use induction and the recurrence relations in (3.32):
@@
(3.35) $\left(\frac{u\left(3t^{2}+du^{2}\right)}{4N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{t\left(t^{2}+3du^{2}\right)}{4N_{\varepsilon}^{2}}\right)y_{k}$ $=-2N_{\varepsilon}y_{k-3}\quad\text{ and }$
$\left(\frac{t\left(t^{2}+3du^{2}\right)}{4N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{du\left(3t^{2}+du^{2}\right)}{4N_{\varepsilon}^{2}}\right)y_{k}$ $=2N_{\varepsilon}x_{k-3},$
@@
Writing $3t^{2}+du^{2}=-4N_{\varepsilon}+4t^{2}$ and $t^{2}+3du^{2}=4N_{\varepsilon}+4du^{2}$, we see that the coefficients of $x_{k}$ and $y_{k}$ in the relationships in (3.35) are integers.
We divide the remainder of the proof into cases according to the parity of $t$ and $u$.
@@
Since $\left(\left(t+u\sqrt{d}\right)/2\right)^{2}=\left(t^{2}+u^{2}d\right)/4+(tu/2)\sqrt{d}=\left(t_{2}+u_{2}\sqrt{d}\right)/2$, where $t^{2}+u^{2}d=\pm 4+2du^{2}\equiv 2\pmod{4}$, it follows that $t_{2}$ and $u_{2}$ are both odd. Hence $x_{2}$ and $y_{2}$ are not integers. However,
$\left(\frac{t+u\sqrt{d}}{2}\right)^{3}=\frac{t^{3}+3tu^{2}d}{8}+\frac{3t^{2}u+u^{3}d}{8}\sqrt{d}=\frac{t_{3}+u_{3}\sqrt{d}}{2},$
where $t^{3}+3tu^{2}d=t\left(t^{2}+3u^{2}d\right)=t\left(\pm 4+4u^{2}d\right)\equiv 0\pmod{8}$ and $3t^{2}u+u^{3}d=-u\left(-4t^{2}\pm 4\right)\equiv 0\pmod{8}$. So both $t_{3}$ and $u_{3}$ are even. Hence $x_{3}$ and $y_{3}$ are both integers.
By an inductive argument using the recurrence relations for the $x_{k}$’s and $y_{k}$’s in (3.32), one finds that $x_{k}$ and $y_{k}$ are both integers if and only if $3|k$, when $t$ and $u$ are both odd and $a$ and $b$ have opposite parities. So we consider only $\gcd\left(x_{3k},y_{3k}\right)$.
Since $t$ and $du$ are both odd in this case, we have $3t^{2}+du^{2}=-4N_{\varepsilon}+4t^{2}\equiv 0\pmod{8}$ and $t^{2}+3du^{2}=4N_{\varepsilon}+4du^{2}\equiv 0\pmod{8}$, so the coefficients of $x_{k}$ and $y_{k}$ in (3.35) can all be divided by $2$ to eliminate the factors of $2$ on the right-hand sides there. Thus, $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-3},y_{k-3}\right)|\cdots|\gcd\left(x_{0},y_{0}\right)=1$.
(ii-b) If $a$ and $b$ have the same parity, then $y_{1}=(au+bt)/2\in\mathbb{Z}$. Similarly, $x_{1}\in\mathbb{Z}$. Using the recurrence relations in (3.32), the $x_{k}$’s and $y_{k}$’s are integers.
As in case (ii-a), we can remove the factor of $2$ on the right-hand side of (3.35). So we find that if $k\equiv k_{1}\pmod{3}$ with $0\leq k_{1}\leq 2$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k_{1}},y_{k_{1}}\right)$.
@@
Suppose first that $y_{0}$ is odd. Since $t$ is odd, we find that $y_{2}=ty_{1}\pm y_{0}$ is even if $y_{1}$ is odd and $y_{2}$ is odd otherwise.
In the first case (when $y_{1}$ is odd), we have $\gcd\left(x_{1},y_{1}\right)|\gcd\left(x_{0},y_{0}\right)$ and hence
@@
Now we suppose that $x_{0}$ is odd. Since $t$ is odd, we find that $x_{2}=tx_{1}+x_{0}$ is even if $x_{1}$ is odd and $x_{2}$ is odd otherwise. As above, we find that
@@
(iii) Now suppose that $t$ is even and $u$ is odd. In this case, $4|d$.
Since $y_{1}=(au+bt)/2$, we see that $y_{1}\in\mathbb{Z}$ if and only if $a$ is even. Hence all the $y_{k}$’s are integers if $a$ is even. Since $t$ is even, using the recurrence relation for the $y_{k}$’s in (3.32), we see that for $a$ odd, $y_{k}\in\mathbb{Z}$ when $k$ is even and $y_{k}\not\in\mathbb{Z}$ otherwise.
Since $x_{1}=(at+bdu)/2$ and $t$ and $d$ are both even, we see that $x_{1}\in\mathbb{Z}$. Hence, by the recurrence relation for the $x_{k}$’s in (3.32), $x_{k}\in\mathbb{Z}$ always holds.
Here we use (3.34), observing that the coefficients of $x_{k}$ and $y_{k}$ on the left-hand sides are all even. So we have relationships with integer coefficients and without the factor of $2$ on the right-hand sides. Hence
$\gcd\left(x_{k},y_{k}\right)\text{ divides }\left\{\begin{array}[]{ll}\gcd\left(x_{0},y_{0}\right)=1&\text{if $k$ is even},\\ \gcd\left(x_{1},y_{1}\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right)=2&\text{if $k$ is odd}.\end{array}\right.$
@@
Let the sequences $\left(x_{k}\right)_{k=-\infty}^{\infty}$ and $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be as defined in (1.1), with the notation and assumptions there. Suppose that $k\neq 0$, $\gcd(a,b)=1$, $N_{\alpha}<0$ and that $x_{k}$ and $y_{k}$ are both integers. Using the notation of Subsection 2.1 with $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$, $u_{1}=2x_{k}$ and $u_{2}=\pm 2\sqrt{N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)}$, then
@@
Put $v_{p}\left(g^{2}\right)=\ell\geq 0$. We will prove that $v_{p}\left(\gcd\left(a^{2},d\right)\right)=\ell$. We also put $v_{p}\left(g_{1}^{2}\right)=\ell_{1}$ and $v_{p}\left(g_{2}\right)=\ell_{2}$. Note that $\ell_{2}=0$ or $1$, since $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$ (from the definition of $g_{2}$, we have $g_{2}|t^{\prime}$ and here $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$). From the definition of $g_{3}$, we have $g_{3}|4$, so $v_{p}\left(g_{3}\right)=0$ for odd primes, $p$. Hence $\ell=\ell_{1}+\ell_{2}$.
@@
If $p\mid\operatorname{sf}\left(N_{\alpha}\right)$, then $p^{\ell_{1}+1}\parallel N_{\alpha}$. We must have $\ell_{2}=0$, since otherwise $\ell_{1}+1<\ell_{1}$ and we get a contradiction to $p^{\ell+1}|N_{\alpha}$. In this case, $p\nmid\left(x_{k}^{2}/g_{1}^{2}\right)$. So $p^{\ell_{1}}\parallel x_{k}^{2}$. But this contradicts $p^{\ell+1}|x_{k}^{2}$. Hence $p^{\ell+1}\nmid\gcd\left(a^{2},d\right)$. Therefore, for all primes $p>2$, $v_{p}\left(g^{2}/\gcd\left(a^{2},d\right)\right)=0$.
@@
We have $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}\right)\right)\geq\ell+2$. Also $v_{2}\left(g_{2}\right)=1$. Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}/g_{3}\right)\geq\ell$, recalling that $g_{3}=1,2$ or $4$.
(2.2) Suppose that $2|\operatorname{sf}\left(N_{\alpha}\right)$.
Here $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(\gcd\left(4x_{k}^{2},4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}/2\right)\right)\geq\ell+1$. We break this case into subcases.
(2.2.1) If $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=\ell$, then $v_{2}\left(N_{\alpha}\right)\geq\ell+1$. So $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}/2\right)\right)\geq\ell+2$. Here $v_{2}\left(g_{2}\right)\geq 1$, so as in case (2.1), we have $v_{2}\left(g^{2}\right)\geq\ell$, as required.
(2.2.2) If $v_{2}\left(a^{2}\right)\neq v_{2}\left(db^{4}\right)$, then $v_{2}\left(N_{\alpha}\right)=\ell$. So $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)=\ell+1$, since $2|\operatorname{sf}\left(N_{\alpha}\right)$. From $v_{2}\left(g_{1}^{2}\right)=\ell+1$, it also follows that $\ell$ must be odd.
From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$ and $v_{2}\left(N_{\alpha}\right)=\ell$, we have either $v_{2}\left(x_{k}^{2}\right)=\ell$ and $v_{2}\left(dy_{k}^{2}\right)\geq\ell+1$; or else $v_{2}\left(x_{k}^{2}\right)\geq\ell+1$ and $v_{2}\left(dy_{k}^{2}\right)=\ell$. The first case is not possible because $\ell$ is odd. In the latter case, we must have $v_{2}(d)=\ell$ and $2\nmid y_{k}$. Also,
$v_{2}\left(u_{1}/g_{1}\right)=(1/2)v_{2}\left(4x_{k}^{2}/g_{1}^{2}\right)\geq(1/2)\left(\ell+3-(\ell+1)\right)=1,$
so $v_{2}\left(g_{2}\right)=1$. Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}g_{2}/g_{3}\right)=v_{2}\left(g_{1}^{2}g_{2}/4\right)=\ell$. ∎
@@
To establish the equality in equation (3.36), we use Lemma 3.11. From that lemma, we can write $g^{2}=2^{m}\gcd\left(a^{2},db^{4}\right)$, where $m\geq 0$. From (2.1) with $n=4$, we find that $\mathcal{N}_{d^{\prime},4}^{2}=2^{\min\left(v_{2}(d^{\prime}),6\right)}$. So, squaring both sides of the equality in (3.36), it is equivalent to
$2^{\min\left(6+m,v_{2}(d^{\prime})+m\right)}=2^{\min\left(6,v_{2}(4b^{\prime})\right)}.$
From the definition of $d^{\prime}$ and Lemma 3.11, we have $d^{\prime}=4N_{\alpha}/g^{2}=4N_{\alpha}/\left(2^{m}\gcd\left(a^{2},db^{4}\right)\right)$ Applying this, along with the definition of $b^{\prime}$, it follows that the equality in (3.36) is equivalent to
@@
To show that this equality holds, we prove that if $v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)>6$, then $m=0$ (i.e., $g^{2}=\gcd\left(a^{2},db^{4}\right)$). Putting $\ell=v_{2}\left(\gcd\left(a^{2},db^{4}\right)\right)$, we must have $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=v_{2}\left(d\right)=\ell$. The second-last equality holds because $\gcd(a,b)=1$. From Lemma 3.10, we have $\gcd\left(x_{k},y_{k}\right)=1,2$. We will show that $y_{k}$ is odd and hence $\gcd\left(x_{k},y_{k}\right)=2$ is not possible here.
We assume that $y_{k}$ is even and derive a contradiction. From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $\left(x_{k}/2^{\ell/2+1}\right)^{2}-\left(d/2^{\ell}\right)\left(y_{k}/2\right)^{2}=N_{\alpha}/2^{\ell+2}$. Since $v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)>6$, we have $v_{2}\left(N_{\alpha}\right)\geq\ell+5$. It follows that $x_{k}/2^{\ell/2+1}$ and $y_{k}/2$ are both odd. Arguing mod $8$, we find that $d\equiv 1\pmod{8}$ and that $t$ and $u$ must both be even.
From (1.2), we can write $4y_{k}=\left(b^{2}\left(2du^{2}\pm 4\right)\pm 2atu\right)$ for some positive integers $t$ and $u$ with $t^{2}-du^{2}=\pm 4$. Since $t$ and $u$ are both even, it follows that $v_{2}\left(4y_{k}\right)=v_{2}\left(2du^{2}\pm 4\right)=2$. This contradicts the assumption that $y_{k}$ is even.
From $y_{k}$ is odd and $\left(x_{k}/2^{\ell/2}\right)^{2}-\left(d/2^{\ell}\right)y_{k}^{2}=N_{\alpha}/2^{\ell}$, we have
$v_{2}\left(g_{1}^{2}\right)=\min\left(v_{2}\left(4x_{k}^{2}\right),v_{2}\left(4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=v_{2}\left(4x_{k}^{2}\right)=\ell+2.$
@@
Suppose that $b=1$, $x_{k}$ and $y_{k}$ are both integers with $y_{k}>1$ and that $g$ and $\mathcal{N}_{d^{\prime},4}$ are as above.
@@
Remark.
Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, $k=-1$, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$.
@@
We have $|g|\mathcal{N}_{d^{\prime},4}\geq 2^{1+\min\left(v_{2}(N_{\alpha})/2,2\right)}\geq 2$ from (3.36). Since $b=1$, from Lemma 3.5(c), we find that $y_{k}\geq\left|N_{\alpha}\right|/4$ holds, so